OpenStudy (anonymous):

Please Help!! There are 50 students total. Out of those 50, there are 18 taking English, 23 are taking Math, and 10 are taking English and Math (the others are taking other subjects so don't worry about them). If one is selected at random, find the probability that the student is taking English or Math. Select one: a. 18/50 b. 23/50 c. 10/50 d. 31/50

OpenStudy (misssunshinexxoxo):

Which one do you believe it is?

OpenStudy (misssunshinexxoxo):

If you don't know that's ok I'll help guide you

OpenStudy (anonymous):

31/50

OpenStudy (misssunshinexxoxo):

what's 18 plus 23?

OpenStudy (misssunshinexxoxo):

18+23= 41 41-10 = 31 31/50 = 0.62 = 62 percent chance

OpenStudy (misssunshinexxoxo):

Correct

OpenStudy (anonymous):

yeah that makes more sense than what my teacher gives us. Thanks

OpenStudy (misssunshinexxoxo):

Cheers:) my pleasure